OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (anonymous):

Point slope form is y-y1=m(x-x1) i think it could be y-2=m(x+3) I could be wrong though

OpenStudy (anonymous):

i think m is -1/5